8-pack control system working group meeting

14 January 2002

In attendence: Chestnut, Frisch, Hall, Kim, Nelson, Rock, Ross

Notes by: J Nelson


(Janice consolidated the action items from all previous meetings, and they are included below as well.)



-Progress from Kukhee: All PWR PVs read out numbers now.  He’s working on implementing the PMAX and PMIN PVs next.

-Kukhee and Ron will work on the database display for PWR ADC channels

-Janice will talk to Steve about an old action item concerning the load out versus hybrid out naming convention

-Janice will work on adding to the panels diagnositc plot buttons for the powers



-No new progress.  They’re done we think.

-Ron will find out if we have enough DACs for 4 klystrons


DAC/ADC Issues

-Janice will work with NLCTA OPS and LLRF to determine if we want phase wraparound and what it should look like.

-Do we need a gold phase for each klystron/time step (SS)?

-the SLED variables need to trickle down to the I/Q and A/P PVs and change their values (SWE, later)



-Progress: SWE got a channel assignment list but still need to make sure the voltages are correct and that all the inputs and outputs are arranged correctly (RC/AY)

-LLRF + NLCTA Ops need to sort out what the LLRF enable and the mod off buttons do


SCP Configs

-All ready with potentially the following exceptions which Janice will check on:

do gold configs exist for the needed areas? & are there buttons for all four klystrons?



-Ron and Kukhee will check to make sure all the waveforms needed exist in the database (consult Steph’s LLRF PV list on the web)

-Joe has one of the modules otherwise they’re all out in the crate

-longer term issue: what to do if reading out an SIS when a fault occurs?

-Right now the SISes are paused and read out with a user request only.  A read when a fault happens will also be added to this.




-Janice will send a list of waveforms to be archived to Bob et al

-SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when a save is requested by the user

-Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)

Other analogs

-Lee Ann got the list from Judy and told Janice the values were being saved to Oracle

-Bob and Lee Ann with switchover NLCTA/8-pack to Oracle sometime this week



-Kukhee made the necessary changes given that the VSTEP was wired in reverse of the original spec.



Janice will ask Steve and Andrew about the 6 calibration variables for each channel as well as (for her own information) what do the little select buttons on the module panel do


Processing algorithm

-Ron and Kukhee reported that the work is ongoing


Peak Power Meter (PPM)

-First we need names for all four PPM channels from Steve then

-We need database for the usual PPM PVs (NLCTA list) plus a waveform for each channel

-Tonee will tweak the labview program to send the waveform to epics once we have names

-Tonee and Janice will set up the necessary hardware in ESB (Enet box, cables, etc)

-Janice will make panels and dm2k plots for these data



-We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archiver stuff as needed



-Judy will make sure crate profile on web is up to date

-Janice will continue to work with Bob Fuller for a dead man switch for epics

-Janice will work with NLCTA Ops to sort out if we need to read out the bucking coil current different from other analogs

-public/private matlab issues to be resolved at a later date

-LLRF needs to decide what the reset button does.